one half of Terri's age is the same as one fourth of Brian's. The sum of their ages is 36. how old is each?
Answer by stanbon(75887)

one half of Terri's age is the same as one fourth of Brian's. The sum of their ages is 36. how old is each?
-----
Equations:
(T/2) = B/4
T + B = 36
---
Rearrange:
T = B/2
----
Substitute for "T" and solve for "B":
B/2 + B = 36
(3/2)B = 36
B = (2/3)(36) = 24 yrs (Brian's age now)
----
solve for "T":
T = B/2
T = 12 (Terri's age now)
============================
Cheers,
Stan H.
=========